the sum of three numbers is 237. the second number is 9 less than two times the first number. the third number is 6 more than three times the first number. what are the three numbers?
Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

x + y + z = 237
.
y = 2x -9
.
z = 3x +6
.
substitute
.
x +2x-9 +3x+6 = 237
.
6x -3 = 237
.
6x = 240
.
x = 40
.
substitute
.
y = 2x -9
y = 2*40 -9
y = 71
.
z = 3x +6
z = 3*40 +6
z = 126
.
Check the sum
.
40 + 71 + 126 = 237
.
Answer: The three numbers are 40, 71, and 126.
.
Done.

Answer by stanbon(57347)

the sum of three numbers is 237. the second number is 9 less than two times the first number. the third number is 6 more than three times the first number. what are the three numbers?
------
Equations:
x + y + z = 237
y = 2x-9
z = 3x+6
---------------
Substitute for "y" and for "x"; then solve for "x":
x + (2x-9) + (3x+6) = 237
---
6x - 3 = 237
6x = 240
x = 40
-----
y = 2x-9 = 80-9 = 71
-----
z = 3x+6 = 126
=======================
Cheers,
Stan H.
=======================